OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 −9 + 2y −9 − 2y 9 − 2y 9 + 2y

OpenStudy (anonymous):

−9 + 2y

OpenStudy (immings):

(times by -4) 4x - 8y = -36 l (times 5) -5x + 10y = 45 minus 4x +5y = 9 l (times 2 then subtract.) (-) 8x + 10y = 18 0x -13y = -45 l -13x = 27 y= -45/-13 l x= 27/-13